as a fish floats on a surface of water its angle of vision is alpha , as it dives into water it becomes beta..... compare the two angles


1
Expert's answer
2021-10-02T14:55:09-0400

"\frac{\sin\alpha}{\sin\beta}=\frac{n_2}{n_1}=n>1,\implies"

"\sin\alpha>\sin\beta,\implies"

"\alpha>\beta."